OpenStudy (anonymous):

find the general form of the equation of a line if it passes through the points a) (1,2) and (3,4) b) (1,3) and (2,5) c) (-1,2) and (-3,-5) d) (-2,4) and (3,4) e) (0,0) and (5,8) f) (2,2) and (5,5)

OpenStudy (anonymous):

it like 3 in the morning and i have another booklet and 2 pages to go please help

OpenStudy (anonymous):

if i dont hand it in by 1 im dead

OpenStudy (midhun.madhu1987):

if the points are (x1 , x2) & (y1 & y2) then general form is : (y-y1)/(y2-y1) = (x-x1)/(x2-x1)

OpenStudy (midhun.madhu1987):

i think so

OpenStudy (lochana):

ok. I will solve first one for you. and you have to find other solutions using the way I solved. so the points are (1,2) and (3,4) these points must be on the solution to find the solution I will use this method \[\frac{ y -2 }{ x - 1 } = \frac{ 2 - 4 }{ 1 -3 }\]

OpenStudy (lochana):

by simplifying you will get y = x +1